OpenStudy (littlenugget):

Solve by extracting the square roots: 3(3x-1)^2=21

OpenStudy (anonymous):

divide by 3 to get \((3x-1)^2=7\)

OpenStudy (anonymous):

then take the square roots

OpenStudy (littlenugget):

3x-1=\[\sqrt{7}\]? Like this?

OpenStudy (anonymous):

yeah don't forget \[3x-1=-\sqrt7\]

OpenStudy (littlenugget):

how come it's negative?

OpenStudy (anonymous):

could be \(\sqrt7\) or could be \(-\sqrt7\)

OpenStudy (littlenugget):

ohhhh ok :) so would that be the answer? or is there more to it?

OpenStudy (anonymous):

yeah you have to solve for \(x\)

OpenStudy (anonymous):

add 1 divide by 3

OpenStudy (littlenugget):

So it would be:\[\frac{ 1\pm \sqrt{7} }{ 3 }\]

OpenStudy (anonymous):

yes

OpenStudy (littlenugget):

so is that what they mean by extracting the square roots?

OpenStudy (anonymous):

i guess

OpenStudy (littlenugget):

okie dokie :) thank you so much!!!!! :D

OpenStudy (anonymous):

yw